I need help with this particular problem, I keep getting y=2/5x-1 but it is not correct, could someone give me a detailed way to solve this problem??
Problem : The equation of a line that goes through the point (8,6) and is parallel to the line 2x+5y=5.
Any help with this will be greatly appreciated.

Answer by Alan3354(69443)

I need help with this particular problem, I keep getting y=2/5x-1 but it is not correct, could someone give me a detailed way to solve this problem??
Problem : The equation of a line that goes through the point (8,6) and is parallel to the line 2x+5y=5.
---------------------------
Find the slope of the line. Put it in slope-intercept form, which means solve for y.
2x+5y=5
5y = -2x + 5
y = (-2/5)x + 5 This is the slope-intercept form, y = mx + b. m is the slope.
m = -2/5
All lines parallel will have the same slope.
then use:
y-y1 = m*(x-x1) where (x1,y1) is the point
y-6 = (-2/5)*(x-8)
y-6 = (-2/5)x + 16/5
y = (-2/5)x + 46/5 slope-intercept form
-----------------------
2x + 5y = 46 standard form